# How Haag-tied is QFT, really?

### 18 Dec 2023 v2

@@ 6 Conclusion
We have argued that these extra-Haagian outlooks are the primary driving forces in the various debates on Haag’s theorem, and that these furthermore shape the different viewpoints as to what are the most important future projects regarding QFT, as well as on the appropriate methods for accomplishing those projects. We have urged stakeholders in the future of QFT to reconsider their own expectations for QFT, reflecting seriously on meta-level questions regarding the nature of foundational work in physics. What role does (should) foundational work play in the progress in physics? How does (should) foundational work coordinate with non-foundational work? What does (should) foundational work even look like? In calling attention to these questions, we have argued that stakeholders in physics, mathematics, and philosophy ought to give more specific characterizations of their proposed methods for making progress, together with more detailed arguments as to why those preferred methods are likely to succeed.

### 06 Apr 2024 v3

@@ 6 Conclusion
We have argued that these extra-Haagian outlooks are the primary driving forces in the various debates on Haag’s theorem, and that these furthermore shape the different viewpoints as to what are the most important future projects regarding QFT, as well as on the appropriate methods for accomplishing those projects. We have urged stakeholders in the future of QFT to reconsider their own expectations for QFT, reflecting seriously on meta-level questions regarding the nature of foundational work in physics. What role does (should) foundational work play in the progress in physics? How does (should) foundational work coordinate with non-foundational work? What does (should) foundational work even look like? In calling attention to these questions, we encourage stakeholders—especially philosophers—to assess the significance of the reasoning surrounding theorems like Haag’s rather than the theorems themselves. This, we believe, promises to deliver more useful guidance by engaging with the reasoning physicists and mathematicians are already doing as they make difficult choices about how to proceed with the construction of QFT.